1.five times the second of three consecutive odd integers is thirteen less than three times the sum of the first and third integers Find the integers

Answer by richwmiller(17219)

some ground rules:
one problem at a time
no similar problems
limit 4 problems a day
5y=-13+3(x+z)
y=x+2
z=x+4